Solve the following problems by defining the variable, writing a quadratic equation, solving it and then providing a written conclusion.
- The sum of the squares of two consecutive even numbers is 1460. Find the numbers.

Answer by stanbon(75887)

The sum of the squares of two consecutive even numbers is 1460. Find the numbers.
--------------------
1st even: 2x
2nd even: 2x+2
--------------------
(2x)^2 + (2x+2)^2 = 1460
4x^2 + 4x^2 + 8x + 4 = 1460
8x^2 + 8x -1456 = 0
x^2 + x - 182 = 0
(x-13)(x+14) = 0
--------------------------
x = 13
-------------
1st: 2x = 26
2nd: 28
=====================
Cheers,
Stan H.